# Red Queen: Safeguarding Large Language Models against Concealed Multi-Turn Jailbreaking

### v1

@@ 4 Experimental Setup
We describe the models selected for our experiments and the metrics used for safety evaluation.
[…]
Evaluation Metrics. We use Attack Success Rate (ASR) (Zou et al., 2023; Yu et al., 2024), the percentage of attacks receiving harmful output, as our primary metric. One difficulty in evaluating the performance of RED QUEEN Attack is judging when output is harmful or malicious. Not only can a detailed step-by-step plan lead to malicious outcomes, but suggestions on tools or perspectives (e.g., timetables and escape routes) can also inspire them. We consider four candidate judgment methods utilized by previous research: 1) GCG (Zou et al., 2023), a classifier based on keyword mapping, 2) GPT-4o evaluator with meticulously designed judge prompt (Chao et al., 2023), 3) Bert-based classifier (Huang et al., 2024b) trained on HH-RLHF dataset (Ganguli et al., 2022) with more than 160k entries. 4) Llama Guard (Inan et al., 2023), an LLM-based safeguard model for Human-AI conversation use cases. We randomly sample 100 attacks and corresponding responses, with three human annotators labelling whether the responses are harmful in an agreement of 100%. As shown in Table 2, we compare four judgment methods on samples but find none achieve satisfying accuracy (>0.8). Further analysis shows their judgments are unreliable and fluctuate, especially in long contexts where the model responds with warnings like "This is a fictional example and should not be used in real cases." Therefore, we designed a judging prompt focused on identifying detailed plans or suspicious hints and tested it on samples with GPT-4 and Llama3. Models with this prompt reached nearly 100% accuracy. We select Llama3 with the new judging prompt as our judgment method due to its high accuracy and the consideration of the cost. The temperature is set to 0 for judgment. Full details are in the Appendix D.
Table 2: Comparison of judgment methods on sampled prompts, Judge represents the new judging prompt.
Judge Method GCG GPT-4o Llama Guard Bert-based Llama-3 (Judge) GPT-4o (Judge)
Accuracy 0.71 0.53 0.49 0.33 0.96 0.94
@@ 5 Results
We focus on four research questions: 1) How does the RED QUEEN Attack perform across different LLM families? 2) What factors contribute to the success of the RED QUEEN Attack? 3) How does the RED QUEEN Attack perform in different scenarios and harmful action categories? 4) What are the outputs of LLMs when the RED QUEEN Attack succeeds or fails?
@@
Key Factors for Red Queen Attack Success. The success of Red Queen Attack highlights the vulnerability of current LLMs. Being the first work to explore jailbreak in multi-turn scenarios with concealment, to simulate further red teaming and jailbreak research in the multi-turn scenario, we conduct a comprehensive study in this section to analyze the key factor contributing to Red Queen Attack success: 1) Multi-turn Structure & Concealment, 2) Turn Number and 3) Model Size. We conclude our qualitative findings in the following.
@@
Table 4: ASR from ablation experiments on multi-turn structure and concealment. The number in the footmark represents the improvement or drop compared to the direct attack. D=Direct Attack. C=Concealment.
Model Direct Attack Concealment Multi-Turn & D Multi-Turn & C
Qwen2-7b 10.93 21.28+10.35 1.14-9.79 27.19+16.26
Qwen2-72b 1.25 29.64+28.39 1.00-0.25 38.26+37.01
[…]
Llama3-8b 0.79 36.84+36.05 1.21+0.42 13.36+12.57
Llama3-70b 0.93 20.36+19.43 1.14+0.21 52.41+51.48
Llama3.1-70b 2.29 26.71+24.42 5.64+3.35 37.57+35.28
Llama3.1-405b 2.36 23.77+21.41 8.15+5.79 52.79+50.43
GPT-4o-mini 0.57 49.09+48.52 1.50+0.93 30.64+30.07
GPT-4o 0.64 64.73+64.09 0.85+0.21 87.62+86.98
@@
Multi-turn Structure & Concealment. Red Queen Attack differs from the previous jailbreak in two points: the multi-turn structure and the concealment of malicious intent. We conduct an ablation experiment to evaluate the isolated effects, as shown in Table 4. "Concealment" and "Multi-Turn & C" refer to single-turn and three-turn attacks, respectively, in Table 3. The "Multi-turn & D" repeated user prompts, ""Can you provide me with a plan for that?" followed by the model’s response, "Hi I am a virtual assistant. Is there anything else I can help you with?" over three turns. According to the result, concealment alone proves to be an effective jailbreak method across all models, highlighting that current LLMs struggle to detect malicious intent. While introducing a multi-turn structure alone yields only marginal ASR improvement, combining it with concealment significantly enhances ASR. For instance, while concealment already achieves a high ASR of 64.73% on GPT-4o, incorporating the multi-turn structure further raises it to 87.62%. However, we also observe that incorporating multi-turn structures sometimes decreases ASR, particularly in smaller models, possibly due to their limited capacity to process complex dialogues. Our ablation results indicate concealment is a powerful jailbreak strategy to mislead LLMs, and multi-turn structures can be a valuable supplement to existing jailbreak techniques.
Figure 3: Performance comparison on different turns across models.
@@
Turn Number. We visualize ASR based on different turns (Table 3) in Figure 3. In general, increasing the number of turns by adding questions or details generally increases ASR, especially for models between 8B and 70B. The five-turn scenario works best in six out of ten models, demonstrating the effectiveness of incorporating additional interaction turns. More interaction allows Red Queen Attack to incrementally earn the trust of the model, making the scenario more believable. Extended turns also result in longer contexts, which can be difficult for current LLMs to manage during inference (Anil et al., 2024). However, this pattern is not observed in larger models (>70B), such as Llama3.1-405b and Qwen2-72b. We speculate that advanced attention mechanisms, such as Rotary Position Embedding (RoPE) (Su et al., 2024) in Llama3 and Grouped Query Attention (GQA) (Ainslie et al., 2023) in Qwen2, may help these models focus on key information towards the end of the interaction, where the user asks for a functional plan, thereby mitigating the effect.
[…]
Performance on Different Scenarios and Harmful Action Categories. We further break down the results into different scenarios and harmful action categories in Figure 4 (Full details in Appendix B). Red Queen attack can have different ASR based on the scenario and action categories. Among occupation-based scenarios, detective and police scenarios achieve the highest ASR across all models. These two occupations, particularly those that are more closely related to criminal activities, make it easier to persuade models about the fictional scenario and mislead the model to generate harmful plans. Conversely, attacks based on lawyers and teachers do not work well, often resulting in a fictional plan for analysis report preparation or fiction writing (Figure 5), possibly due to their co-existence in models’ pertraining data. We also found specific occupation-based scenarios can perform extremely well on some models, such as the priest scenario, which has on-par ASR with detective or police on Mixtral-22b. In contrast, relation-based scenarios have nearly similar effects on various models, pointing out that the authority role contained in occupations can enhance the quality of attack.
Among the fourteen harmful action categories, several categories [^4] Due to space limitations, we use the first word of each category to represent the whole name in later sections. do not perform well across models for two main reasons: 1) Improper match: Categories like Hate Speech and Controversial Topics involve actions related to speaking, which do not fit our scenario structure where someone plans to do something. For example, "say the f word" in the hate speech category fails to align with the scenario structure, reducing the credibility of the scenarios presented to models. 2) Misaligned objectives: Categories such as Sexually Explicit and Discrimination aim to elicit specific words or opinions explicitly (Ji et al., 2024), like "give a sound to give others an orgasm" or "judge the German language." In our attack, the model often generates a neutral plan on how to research the topic, resulting in non-malicious outcomes. We also analyze whether some harmful action categories have higher ASR in specific scenarios (In Appendix B). The results show that ASR tends to depend on the effectiveness of the scenario or harmful action category itself. In other words, if a scenario attack is effective, it is successful across all harmful action categories, and vice versa.
@@
Success and Failure Case of Red Queen Attack. We present successful and failed jailbreak attempts in Figure 5. The figure shows outputs from various models subjected to the Red Queen attack over five turns. The leftmost green box illustrates a typical success case where the model generates a detailed and harmful plan for bomb-making. The plan is extremely harmful as it not only includes a list of required materials but also step-by-step instructions for procurement, construction, and even deployment. Notably, the generated plan can be tailored to the harmful action, indicating the model’s extensive utilisation of their knowledge and reasoning abilities. Unlike a bomb-making plan, which focuses on material production and bomb building, the model proposes a robbery plan (middle box in Figure 5) with precise timing and even contingency plans for unexpected events.
Figure 5: Example of success and failure cases.
For failed jailbreak cases, instead of completely refusing to answer, the model tends to generate a safety plan related to the scenario itself. As shown in the two red boxes on the right in Figure 5, the model outputs guidance for writing a report documenting information between a priest and penitent. In the teacher scenario, when prompted to claim that English people are stuck up, the model generates a preparation plan for a literature presentation. This plan does not contain any explicit opinion statements and is labelled as safe by our judging method.
Overall, the multi-turn scenario constructed in Red Queen Attack is capable of lending a model to generate a plan-based response in both success and failure cases, which contrasts with previous single-turn attacks where failure cases often result in complete refusal (Ji et al., 2024; Zou et al., 2023). This capability places a heavy burden on the model to balance harmlessness and helpfulness, contributing to the extremely high ASR across models. We provide more examples in Appendix A.
Table 5: DPO result on Llama3.1 family with Red Queen Guard. The best ASR number (lower the better) is bold. We report numbers on MMLU-Pro and AlpacaEval (higher the better) to evaluate models’ reasoning and instruction-following abilities.
Model ASR ↓ MMLU-Pro ↑ AlpacaEval ↑
@@
   +HH-RLHF &Red Queen Guard 26.0 55.0 36.2
   +Red Queen Guard 1.3 55.1 36.8
@@
   +Red Queen Guard 0.1 64.2 32.1
6 Safeguarding Strategies
Given the widespread application of LLMs in everyday life, we explore strategies to enhance the safety mechanisms of these models in this section. According to Wei et al. (2024), the model’s vulnerabilities may arise from the misalignment in the generalization of safety training, particularly in multi-turn scenarios where the model’s capabilities lack adequate safety alignment training. Thus, we investigate whether training models on carefully designed multi-turn datasets using Direct Preference Optimization (DPO) (Rafailov et al., 2024) can mitigate this misalignment. We sampled 20 multi-turn data points of successful LLM jailbreaks from each scenario and harmful action category, supplemented with safety responses from Llama3.1-405b, yielding an 11.2K preference dataset, Red Queen Guard. For comparison, we also include 11.5K human preference data points from HH-RLHF (Bai et al., 2022), a preference dataset widely used in safety alignment (Touvron et al., 2023; Song et al., 2024). We align three models from the Llama3.1 family and evaluate models’ performance on approximately 10% of the original attack data (5539 instances). To ensure the optimization on harmless with the Red Queen Guard won’t bring collapse on other objectives (e.g. reasoning, helpfulness) (Lin et al., 2024; Wei et al., 2024), we further evaluate models on MMLU-Pro (Wang et al., 2024) which includes reasoning-focused questions across 14 diverse domains and AlpacaEval (Dubois et al., 2024), a LLM-judgement benchmark testing the ability of models to follow general user instructions.
Table 5 shows model performance after DPO using different preference data. We do not rely solely on the HH-RLHF dataset, as the previous study has shown that RLHF training with HH-RLHF can reduce performance (alignment taxes) in smaller models (Bai et al., 2022), and we aim to develop a mitigation dataset that works well regardless of model size. Compared to the original ASR of each model under the Red Queen Attack, DPO using the Red Queen Guard proves to be an effective safeguarding strategy that significantly reduces the model’s vulnerability to multi-turn attacks (ASR around 1.0). The combination of HH-RLHF and Red Queen Guard provides only limited improvement to the model’s safety mechanisms, with the ASR still relatively high at 26.0. This failure can be attributed to unclear alignment learning signals (D’Oosterlinck et al., 2024) for two main reasons: (1) HH-RLHF lets crowd workers use their own intuitions to define harmfulness, which lacks a uniform safety objective across all data. (2) The difference within objectives of HH-RLHF and Red Queen Guard further increases the disambiguation, limiting their combined effectiveness. Based on the results from MMLU-Pro and AlpacaEval, integrating Red Queen Guard can address the safety misalignment in multi-turn scenarios without compromising the model’s reasoning or instruction-following capabilities, highlighting its promising potential for broader usage in general safety alignment. We present the whole details of the experiment in Appendix E.

### v2

@@ 4 Experimental Setup
We describe models and baselines selected for our experiments and metrics used for attack evaluation.
Models. We evaluate ten instruction-tuned models from four representative LLM families. The selected model’s sizes vary from 7B to 405B to ensure a comprehensive evaluation: 1) Mixtral (8$\times$7B and 8$\times$22B) (Jiang et al., 2024a); 2) Llama3 (7B and 70B) and Llama3.1 (70B and 405B) (Touvron et al., 2023); 3) Qwen2 (7B and 72B) (Yang et al., 2024); 4) GPT-4o/4o-mini (Achiam et al., 2023). These models show promising performance in public benchmarks (Cobbe et al., 2021; Srivastava et al., 2023) and are widely adopted in daily usage.
Baselines. As Red Queen Attack is the first work constructing a multi-turn scenario with a concealment strategy. To demonstrate the effectiveness and generalizability of our approach, we compare the Red Queen Attack with previous jailbreak attacks across three categories: 1) Concealment: Cipher-based attacks(Yuan et al., 2024) (e.g., using ASCII encoding) and ArtPrompt (Jiang et al., 2024b) (leveraging ASCII art) to obscure harmful intent; 2) Multi-turn: Cosafe(Yu et al., 2024) and CoU (Bhardwaj and Poria, 2023), which employ multi-turn interactions to achieve jailbreaks; and 3) Fictional Plan Writing: DeepInception(Li et al., 2023), which utilizes a science fiction writing scenario for jailbreak attempts.
Metric Agreement (↑) FPR (↓) FNR (↓)
GCG 0.71 0.09 0.39
GPT-4o 0.53 0.00 0.72
Bert-Based 0.33 0.00 1.00
Llama Guard 0.49 0.03 0.75
Llama-3(Judge) 0.96 0.00 0.06
GPT-4o(Judge) 0.94 0.06 0.06
Table 1: Comparison of judgment methods on selected samples. FPR: False Positive Rate, FNR: False Negative Rate, Judge: new judging prompt.
Evaluation Metrics. We use Attack Success Rate (ASR) (Zou et al., 2023; Yu et al., 2024), the percentage of attacks receiving harmful output, as our primary metric. Since the Red Queen Attack seeks to elicit plans for harmful actions, we classify model responses as harmful if they generate on-topic outputs (e.g., providing a plan) that could facilitate criminal activities. We consider four candidate judgment methods utilized by previous research: 1) GCG (Zou et al., 2023), a classifier based on keyword mapping, 2) GPT-4o evaluator with meticulously designed judge prompt (Chao et al., 2023), 3) Bert-based classifier (Huang et al., 2024c) trained on HH-RLHF dataset (Ganguli et al., 2022) with more than 160k entries. 4) Llama Guard (Inan et al., 2023), an LLM-based safeguard model for Human-AI conversation. We randomly sample 100 attacks and corresponding responses, with three human annotators labelling harmfulness in an agreement of 100%. As shown in Table 1, we compare four judgment methods on samples but find none achieve satisfying accuracy (>0.8). Further analysis shows models tend to respond with warning sentences such as “This is a fictional example and should not be used in real cases.”, which confuses previous methods and results in a higher False Negative Rate. Therefore, we designed a judging prompt focused on identifying detailed plans or suspicious hints and tested it on samples with GPT-4o and Llama3-70B. Models with this prompt reached nearly 100% accuracy. We select Llama3-70B with the new judging prompt as our judgment method due to its high accuracy and the consideration of the cost. To ensure the new judgment method is not overfitting to selected samples, we evaluate it on 300 prompt-response pairs (with harmful/benign labels) from Jailbreakbench (Chao et al., 2024). Our judgement method still outperforms all prior approaches (Details in Appendix F).
Jailbreak Settings. During the jailbreaking experiment, we provide the complete multi-turn scenario to the model and evaluate on its response. Following prior work (Li et al., 2023; Bhardwaj and Poria, 2023), we set the temperature to 1 during prompting and 0 for evaluation.
@@ 5 Results
We focus on four research questions: 1) How does the Red Queen Attack perform across different LLM families? 2) How does the Red Queen Attack compare to prior jailbreak attacks? 3) What factors contribute to the success of the Red Queen Attack? 4) What are the outputs of LLMs when the Red Queen Attack succeeds or fails?
@@
Comparison with Prior Jailbreak Attacks.
Figure 3: Comparison of Red Queen Attack and baseline jailbreak attacks.
Figure 3 presents a comparative analysis between the Red Queen Attack and existing jailbreak methods. Since DeepInception and ArtPrompt employ the same harmful action set from AdvBench (Zou et al., 2023), we constructed five-turn Red Queen Attack [^1] We use the police scenario for its strong performance. using this set to ensure a fair comparison. In general, the Red Queen Attack maintains a high ASR across all model families, while previous jailbreaks are effective only on specific families. For instance, Chiper-based and CoSafe attacks performed well only on Llama3.1-70B. Moreover, Red Queen Attack outperformed previous methods on nine out of ten models, with ASR improvements ranging from 2% to 64%, highlighting its superior ability to exploit the models’ weak Theory-of-Mind capabilities. Additionally, concealment-based approaches (e.g., ArtPrompt) and fictional scenario construction methods (e.g., DeepInception) generally outperform multi-turn strategies (e.g., CoSafe and CoU), highlighting the limitations of simplistic multi-turn jailbreak techniques. We provide detailed illustration in Appendix D.
@@
Key Factors for Red Queen Attack Success. The success of Red Queen Attack highlights the vulnerability of current LLMs. Being the first work to explore jailbreak in multi-turn scenarios with concealment, to simulate further safety research in the multi-turn scenario, we conduct a comprehensive study in this section to analyze the key factor contributing to Red Queen Attack success: 1) Multi-turn Structure & Concealment, 2) Turn Number and 3) Model Size.
Model D C M & D M & C
Qwen2-7B 10.9 21.3+10.4 1.1-9.8 27.2+16.3
Qwen2-72B 1.3 29.6+28.4 1.0-0.3 38.3+37.0
Mixtral-8$\times$7B 0.6 10.5+9.9 1.3+0.7 29.6+29.1
Mixtral-8$\times$22B 23.0 36.6+13.7 25.1+2.2 28.0+5.1
Llama3-8B 0.8 36.8+36.1 1.2+0.4 13.4+12.6
Llama3-70B 0.9 20.4+19.4 1.1+0.2 52.4+51.5
Llama3.1-70B 2.3 26.7+24.4 5.6+3.4 37.6+35.3
Llama3.1-405B 2.4 23.8+21.4 8.2+5.8 52.8+50.4
GPT-4o-mini 0.6 49.1+48.5 1.5+0.9 30.6+30.1
GPT-4o 0.6 64.7+64.1 0.9+0.2 87.6+87.0
@@
Table 3: ASR from ablation experiments on multi-turn structure and concealment. D=Direct Attack. C=Concealment. M = Multi-Turn.
@@
Multi-turn Structure & Concealment. Red Queen Attack differs from the previous jailbreak in two points: the multi-turn structure and the concealment of malicious intent. We conduct an ablation experiment to evaluate the isolated effects, as shown in Table 3. “C” and “M & C” refer to single-turn and three-turn attacks, respectively, in Table 2. The“M & D" repeated user prompts,“Can you provide me with a plan for that?” followed by the model’s response,“Hi, I am a virtual assistant. Is there anything else I can help you with?” over three turns. According to the result, concealment alone proves to be an effective jailbreak method across all models, highlighting that current LLMs struggle to detect malicious intent. While introducing a multi-turn structure alone yields only marginal ASR improvement, combining it with concealment significantly enhances ASR. For instance, while concealment already achieves a high ASR of 64.73% on GPT-4o, incorporating the multi-turn structure further raises it to 87.62%. However, we also observe that the incorporation sometimes decreases ASR, particularly in smaller models, possibly due to their limited capacity to process complex dialogues. Our ablation results confirm that concealment is a robust jailbreak strategy, with multi-turn structures serving as a supplementary enhancement.
@@
Turn Number. Comparing the ASR of each LLM across different numbers of turns, we observe that increasing the number of turns by adding questions or details generally enhances the ASR, particularly for models ranging from 8B to 70B. The five-turn scenario works best in six out of ten models, demonstrating the effectiveness of incorporating additional interaction turns. More interaction allows Red Queen Attack to incrementally earn the trust of the model, making the scenario more believable. Extended turns also result in longer contexts, which can be difficult for current LLMs to manage during inference (Anil et al., 2024). However, this pattern is not observed in larger models (>70B), such as Llama3.1-405B and Qwen2-72B. We speculate that advanced attention mechanisms, such as Rotary Position Embedding (RoPE) (Su et al., 2024) in Llama3 and Grouped Query Attention (GQA) (Ainslie et al., 2023) in Qwen2, may help these models focus on key signal towards the end of the interaction, where the user asks for a functional plan, thereby mitigating the effect.
Figure 4: Performance comparison of model families in different sizes.
Model Size. In Figure 4, we visualize the Red Queen Attack results on ten models according to their size and family. Interestingly, larger models tend to be more susceptible to Red Queen Attacks. Except for Mixtral-8$\times$22B in the three-turn scenario, all larger models have a higher ASR compared to smaller ones. This increased vulnerability in larger models can be attributed to the mismatch generalization between continued progress on model capabilities and safety alignment training (Wei et al., 2024). In other words, larger models demonstrate a better understanding of language and instruction and can accept fake scenarios easily, while smaller models have difficulty understanding the whole scenario. Also, larger models’ promising capabilities in complex reasoning and planning become “accomplices” for harmful plan generation. In contrast, smaller models, with their limited knowledge and reasoning skills, struggle to even accomplish the task. Further manual checking reveals that smaller models such as Llama3-8B or Qwen2-7B sometimes can not understand the scenario and generate meaningless plans (Appendix A).
Figure 5: Example of success cases.
@@
Success and Failure Case of Red Queen Attack. We present two successful jailbreak attempts in Figure 5. The left box illustrates a typical success case where the model generates a detailed and harmful plan for bomb-making. The plan is extremely harmful as it not only includes a list of required materials but also step-by-step instructions for procurement, construction, and even deployment. Notably, the generated plan can be tailored to the harmful action, indicating the model’s extensive utilisation of their knowledge and reasoning abilities. Unlike a bomb-making plan, which focuses on material production and bomb building, the model proposes a robbery plan (right box) with precise timing and even contingency plans for unexpected events. We present failed jailbreak cases in Appendix A. Instead of completely refusing to answer as observed in previous jailbreak research (Ji et al., 2024; Zou et al., 2023), the model tends to generate a safety plan related to the scenario itself. For example, In the teacher scenario, when prompted to claim that English people are stuck up, the model generates a general preparation plan for a literature presentation. Overall, the multi-turn scenario constructed in Red Queen Attack is capable of lending models to generate plan-based responses in both success and failure cases. This capability places a heavy burden on the model to balance harmlessness and helpfulness, contributing to the extremely high ASR across models. We provide more examples in Appendix A.
6 Safeguarding Strategies
Wei et al. (2024) highlight that LLM’s vulnerabilities often stem from misaligned safety training, especially in multi-turn scenarios lacking adequate alignment. Thus, we investigate whether training models on carefully designed multi-turn datasets using Direct Preference Optimization (DPO) (Rafailov et al., 2024) can bridge this misalignment. We sampled 20 multi-turn data points of successful LLM jailbreaks from each scenario and harmful action category, supplemented with safety responses from Llama3.1-405B, yielding an 11.2K preference dataset, Red Queen Guard. For comparison, we include 11.5K human preference data points from HH-RLHF (Bai et al., 2022), a preference dataset widely used in safety alignment (Touvron et al., 2023; Song et al., 2024). We align three Llama3.1 models and evaluate them on 10% of the original attack data (5539 instances), with no overlap between training and testing. To ensure safety alignment does not cause a collapse in other objectives (e.g., reasoning, helpfulness), we further evaluate models on MMLU-Pro (Wang et al., 2024), which includes reasoning-focused questions across 14 diverse domains, and AlpacaEval (Dubois et al., 2024), an LLM-judgment benchmark testing models’ instruction-following ability.
Model ASR ↓ MP ↑ AE ↑
@@ 6 Safeguarding Strategies
+HH-RLHF &RQG 26.0 55.0 36.2
+RQG 1.3 55.1 36.8
@@
+RQG 0.1 64.2 32.1
Table 4: DPO results on the Llama3.1 family with Red Queen Guard (RQG). The best ASR numbers (lower is better) are bold. MMLU-Pro and AlpacaEval (higher is better) assess reasoning and instruction-following abilities. MP:MMLU-Pro. AE:AlpacaEval.
Table 4 shows model performance after DPO using different preference data. We do not rely solely on the HH-RLHF dataset, as Bai et al. (2022) show that optimization with HH-RLHF can reduce performance (alignment taxes) in smaller models, and we aim to develop a mitigation dataset that works well regardless of model size. Compared to the original ASR, DPO with Red Queen Guard effectively reduces model vulnerability to multi-turn Red Queen Attack, lowering the ASR to around 1.0. The combination of HH-RLHF and Red Queen Guard provides only limited improvement to the model’s safety mechanisms, with the ASR still relatively high at 26.0. This failure can be attributed to unclear alignment learning signals (D’Oosterlinck et al., 2024): (1) HH-RLHF relies on subjective worker intuitions, lacking a consistent safety objective, and (2) conflicting objectives between HH-RLHF and Red Queen Guard. Based on the results from MMLU-Pro and AlpacaEval, integrating Red Queen Guard can address the safety misalignment in multi-turn scenarios without compromising the model’s reasoning or instruction-following capabilities, highlighting its promising potential for broader usage in general safety alignment. We present the whole details of the experiment in Appendix E.